OpenStudy (anonymous):

what is the graph of 6x + 3y=18

OpenStudy (anonymous):

\[6x + 3y=18\]\[ 3y=-6x+18\]\[ y=-2x+6\]

OpenStudy (anonymous):

First, it would help to put the equation in slope-intercept form. That is, y=mx+b

OpenStudy (anonymous):

thanks:) fbi2015 and 1 lorax2

OpenStudy (anonymous):

Your welcome! Is that all you need?